# Small Batch Size Training for Language Models: When Vanilla SGD Works, and Why Gradient Accumulation Is Wasteful

### v1

@@ 4 Experimental Results
Recent works find that larger batch sizes and relatively sophisticated optimizers are necessary for stable language model training. For instance, Zhao et al. [18] and Kunstner et al. [55] observe that SGD has an extremely slow convergence speed for training language models in comparison to adaptive optimizers like Adam. Other works find that Adam performs poorly with small batch sizes [52, 15, 56, 57].
[…]
@@ 4.1 Small Batch Sizes Render Sophisticated Optimizers Unnecessary
We train the 30M model on 600 million tokens from the FineWeb-Edu dataset [60], following Chinchilla scaling laws [21]. Our model adopts modern design choices including rotary embeddings [61], QK-normalization [62], RMSNorm [63], GELU activations [64], and separate input and output embeddings. We use a context length of 512 and tokenize the dataset using the GPT-2 tokenizer. We provide additional experimental details in Appendix A.
@@
Why is momentum less necessary for small batch sizes? Momentum can be seen as a means of dampening oscillations that occur on ill-conditioned loss functions [65]. When we take large steps, we may overshoot the minimum across short (i.e. high curvature) axes, leading to oscillations. By averaging gradients across iterations, these oscillations cancel out in the momentum term. However, when the optimizer takes small steps, as is the case in small batch training paired with a small learning rate, parameter updates do not overshoot the minimum across short axes, so there are no oscillations to damp. We illustrate this effect in Appendix B through a toy example.
@@ 4.2 Large Batch Training is Sensitive to Hyperparameters
To evaluate robustness to hyperparameters at each batch size, we perform an ablation around the optimal hyperparameters achieving the lowest validation loss for the Adam optimizer. We use the same 30M parameter model trained on the FineWeb-Edu dataset as in Section 4.1.
We report in Figure 3 the change in the loss with respect to the lowest value of the loss achieved at every batch size, ranging from $1$ to $4096$. On the x-axis, we report the scaled value of the hyperparameters relative to the value of the hyperparameter that achieved the best loss. We perform this rescaling to make the plots for different batch sizes easier to compare, since the scale of their optimal hyperparameters varies.
Figure 3: Small batch sizes are robust to hyperparameter misspecification. Loss sensitivity to learning rate, $\beta_{1}$, and $\beta_{2}$ using Adam at batch sizes from 1 to 4096. Each curve sweeps a single hyperparameter while holding the others fixed at their optimal values. We observe that smaller batch sizes exhibit broader low-loss regions across all hyperparameters, indicating greater robustness to misspecification. The x-axis shows a scaling of each hyperparameter relative to its optimum value. We parameterize $\beta_{1}$ and $\beta_{2}$ in terms of their decay half-lives, as described in Section 2. Results are shown for the 30M parameter model trained on 600M tokens of FineWeb-Edu.
@@ 4.3 Adam Hyperparameters: How to Scale Them with Batch Size and Why That Is Necessary
We use our grid search results on the 30M model to test existing scaling laws for Adam hyperparameters and also devise new scaling heuristics for these hyperparameters. Figure 4 shows the validation loss on FineWeb-Edu for different Adam hyperparameters, namely the learning rate, and decay parameters $\beta_{1}$ and $\beta_{2}$. We similarly plot the loss for different values of the second moment half-life $t_{2}$, which refers to the number of tokens it takes for a gradient’s contribution to the momentum to be reduced to half of its initial value. We provide a more detailed description of $t_{2}$ in Section 2.
For the learning rate, we observe that the square root scaling recommended by several works and commonly used in practice does not hold [66, 8, 36, 67, 68]. In fact, Figure 4 (left) shows that the learning rate scales more slowly than a square root factor with the batch size. For instance, as we scale the batch size from 1 to 1024, the square root rule would indicate that the corresponding learning rate should be scaled by a factor of 32, whereas we find that a factor of only about 3 empirically works better.
@@
Next, we test our scaling heuristics in a different setting proposed in Xiao [15]. The authors train a decoder-only transformer with 19M non-embedding parameters using Adam on the C4 dataset [54], and only tune the learning rate with $\beta_{1}$ and $\beta_{2}$ fixed to their default values of $0.9$ and $0.95$, respectively. We reproduce their Figure 10(a) results in Figure 5 (left), where small batch sizes of 16 and 32 perform worse than larger batch sizes 128 and 256.
Figure 5: Scaling $\beta_{2}$ according to our heuristics significantly improves performance of small batch sizes. Left: Results from Xiao [15], showing suboptimal performance for small batch sizes due to fixed Adam hyperparameters. Right: Our results after scaling $\beta_{2}$ such that the decay half-life $t_{2}$ is constant across batch sizes. The smallest batch sizes now perform on par with larger batch sizes.
While our scaling heuristics imply that $\beta_{1}=0.9$ could work well across batch sizes, we recommend scaling $\beta_{2}$ such that $t_{2}$ is fixed. By following this simple one-shot scaling recommendation and without any additional hyperparameter tuning, we achieve significantly different results in Figure 5 (right). Strikingly, small batch sizes 16 and 32 achieve on par performance with larger batch sizes 128 and 256, therefore implying that: (i) the previous results from Xiao [15] are largely a consequence of incomplete hyperparameter tuning, (ii) our simple one-shot heuristic generalizes to other settings.
@@ 4.4 SGD Performs Competitively with Adaptive Optimizers at the Billion+ Parameter Scale
Through our experiments with 30M parameter models, we conclude that: (1) small batch sizes can perform on par with or better than larger batch sizes, (2) small batch sizes require less sophisticated optimizers and are more robust to hyperparameter misspecification, (3) in particular, vanilla SGD can be competitive with adaptive optimizers in the small batch regime, and (4) we propose a new scaling heuristic for $\beta_{2}$ such that the half-life $t_{2}$ is fixed, and validate that it leads to improved performance. In Figure 6, we test these findings at a larger scale by training GPT-2 (124M) [59] and GPT-3 (1.3B) [1] on the FineWeb dataset [60].
For the larger 1.3B model, hyperparameter tuning of the learning rate for each optimizer was computationally prohibitive; therefore, we used AdamW with the default hyperparameter settings from Brown et al. [1] (batch size $0.1512,\beta_{1}=0.9,\beta_{2}=0.95,w=$ 0.1) as our baseline. When using batch size 1 with Adam, we do not perform any hyperparameter tuning – we turn off weight decay and scale the learning rate following our empirical results from Section 4.3. We compare two scaling heuristics for $\beta_{2}$: keeping $\beta_{2}$ fixed and keeping the half-life $t_{2}$ fixed. We do, however, sweep over four learning rates for Adafactor and SGD, since we have no reference values for these optimizers. Interestingly, we find that SGD performs on par with the AdamW baseline, and Adam and Adafactor with batch size 1 both outperform the AdamW baseline.
@@
Figure 6: Vanilla SGD performs competitively at larger model scales with minimal tuning. We train GPT-3 (1.3B) and GPT-2 (124M) on FineWeb and compare validation loss across optimizers. Left: For GPT-3 (1.3B), vanilla SGD with batch size 1 and no momentum performs on par with the AdamW configuration from Brown et al. [1], despite having no optimizer state. Our proposed $\beta_{2}$ scaling rule improves Adam’s performance at small batch sizes and outperforms the baseline. Right: After tuning the learning rate of each optimizer on GPT-2 (124M), Adam and Adafactor at batch size 1 match the performance of AdamW at batch size 512, while SGD performs slightly worse.
Figure 7: Memory-efficient fine-tuning. We fine-tune and evaluate a non-instruction-tuned Gemma 3 (4B) model [69] on the MATH dataset [70]. We compare the use of LoRA ($r=16$) against full model fine-tuning, and a small vs. large batch size. We set the width of each bar proportional to the memory footprint required for training. We fix the second moment half-life to be consistent across batch sizes ($\beta_{2}=0.95$ for batch size 16 and $\beta_{2}=0.997$ for batch size 1). We run each optimizer for 5 epochs, using 8 random seeds, and sweep over 8 learning rates. We train LoRA using Adam and batch size 1.
@@ 4.5 Full Fine-tuning on a Budget
In Sections 4.1, 4.2, 4.3 and 4.4, we studied pretraining across different model scales, always training from scratch for a single epoch. In Figure 7, we show that our findings extend to the fine-tuning setting, even when training for multiple epochs.
@@
Because memory can often be a constraint when fine-tuning large models (more on this in Section 5.1), it is a common practice in fine-tuning to freeze the model weights and only train Low-Rank Adaptation (LoRA) [71] modules, which reduces the number of trainable parameters and thus also reduces the required memory for training.
In Figure 7, we observe that Adafactor with batch size 1 performs comparably to Adam, while having a similar memory footprint to LoRA. We also see that Adam with batch size 1 and hyperparameter scaling from Equation 2 performs at least as well as Adam with batch size 16 in the fine-tuning setting too.
LoRA, however, performs noticeably worse than full parameter fine-tuning, which has also been observed in prior works [72, 73]. For this reason, we argue that in a memory-constrained setting, instead of using LoRA, it is likely preferable to perform full fine-tuning with a small batch size, which obviates the need to store accumulated gradients and enables the use of memory-efficient optimizers like Adafactor without losing performance.
5 Practical Recommendations
5.1 Memory Usage
Figure 8: Minimum memory requirements to train GPT-3 (13.0B), assuming all parameters and activations are stored in 16-bit precision and gradient checkpointing is used after every transformer block layer. The dashed line shows the available memory on an NVIDIA A100 40GB GPU, which is not large enough to store expensive optimizer states or accumulated gradients. We measured the true memory usage of each optimizer by performing fused backward passes in PyTorch on an NVIDIA B200 GPU.
Frontier LLMs are typically trained on large clusters of GPUs that necessitate large batch sizes, and, by extension, also sophisticated optimizers. However, when pretraining, and especially fine-tuning, on a single device, memory can easily become a bottleneck. We need to store the model, the optimizer state, the forward-pass activations, and optionally also accumulated gradients in memory.
We illustrate this point in Figure 8 using the larger GPT-3 (13.0B) model as an example. Assume we store all values in bfloat16 [74] precision and checkpoint gradients after every transformer block layer, so that we only need to store $D$ activation values per layer per token [75]. Further assume that the backward pass is fused with the optimizer update step, so that the full gradient never needs to be materialized in memory. Under this setting, attempting to replicate the common practice of using Adam with a large (accumulated) batch size is infeasible on any consumer-level GPU. In contrast, it is completely feasible to train the model on a consumer-level GPU when using SGD, which is stateless, or Adafactor, which has a negligible state size. Crucially, as shown in Sections 4.1, 4.2, 4.3, 4.4 and 4.5, these simple optimizers achieve the best performance with small batch sizes, matching or even outperforming more sophisticated optimizers using large batch sizes.
@@ 5.1 Memory Usage
Summary: We recommend that practitioners training large models in memory-constrained settings exploit the benefits of small batch sizes rather than trying to emulate the large batch size setting (e.g., through gradient accumulation) typically used in industry.
5.2 Hardware Utilization
Our general recommendation for setting the batch size is to use the smallest batch size that maximizes model throughput (measured in tokens / second) or equivalently the model FLOPs utilization (MFU) [76]. To achieve this objective, we must avoid being bottlenecked by memory bandwidth, which in practice means using a batch size of at least several hundred tokens per device [75]. Additionally, second-order optimizers like Muon introduce a fixed computational overhead per optimizer step. If the batch size becomes very small, this overhead can result in reduced model throughput. If any of these constraints become a bottleneck, we recommend increasing the batch size.
6 Discussion
@@
We show that small batch sizes, even as low as 1, consistently match or outperform large batches across model scales when tuned properly, in both pretraining and fine-tuning. In particular, SGD with batch size 1 performs on par with the AdamW baseline for GPT-3 (1.3B) with minimal tuning. We also propose a simple scaling rule for $\beta_{2}$ based on a fixed second-moment half-life in tokens, which generalizes across model scales. Our results contradict common assumptions about large batch training and gradient accumulation. We recommend using the smallest batch size that maximizes throughput, avoiding gradient accumulation, and relying on simpler optimizers when memory is constrained.
@@
Several open questions remain: How do these findings interact with batch size schedules? Can we design more effective optimizers with small state, tuned to the small batch size regime? And can we theoretically understand why fixing second-moment half-life in tokens generalizes across scales?

### v2

@@ 4 Experimental Results
Recent works find that larger batch sizes and relatively sophisticated optimizers are necessary for stable language model training. For instance, Zhao et al. (2025), Kunstner et al. (2024), and Zhang et al. (2024b) observe that SGD has an extremely slow convergence speed for training language models in comparison to adaptive optimizers like Adam. Other works find that Adam performs poorly with small batch sizes (Vyas et al., 2023; Xiao, 2024; Liu and Ma, 2024; Li et al., 2024).
In this section, we revisit and challenge these beliefs by noting that such findings arise due to poor hyperparameter choices and specific experimental setups, for example, not adjusting the decay parameters $\beta_{1}$ and $\beta_{2}$ for Adam at small batch sizes, or using too large a batch size for SGD. We put these beliefs to the test by running a thorough grid search over optimizer hyperparameters and batch sizes for SGD, Adam, Adafactor, and Muon. As the cost of a grid search grows exponentially with the number of hyperparameters, we only perform exhaustive grid searches on a small transformer decoder-only model with 30 million active parameters following Liu et al. (2024) and Xiao (2024). We use these results to propose heuristics for scaling hyperparameters across batch sizes, and we validate our findings at a larger scale by pretraining GPT-2 (124M) (Radford et al., 2019) and GPT-3 (1.3B) (Brown et al., 2020) and fine-tuning Gemma 3 (4B) (Team et al., 2025). In our experiments, we study batch sizes spanning more than four orders of magnitude: $\{1,4,16,64,256,1024,\text{ and }4096\}$.
@@ 4.1 Small Batch Sizes Render Sophisticated Optimizers Unnecessary
We train the 30M model on 600 million tokens from the FineWeb-Edu dataset (Penedo et al., 2024), following Chinchilla scaling laws (Hoffmann et al., 2022). Our model adopts modern design choice, including rotary embeddings (Su et al., 2024), QK-normalization Dehghani et al. (2023), RMSNorm (Zhang and Sennrich, 2019), GELU activations (Hendrycks and Gimpel, 2016), and separate input and output embeddings. We use a context length of 512 and tokenize the dataset using the GPT-2 tokenizer. We provide additional experimental details in Appendix˜A.
@@
Why is momentum less necessary for small batch sizes? Momentum can be seen as a means of dampening oscillations that occur on ill-conditioned loss functions (Goh, 2017). When we take large steps, we may overshoot the minimum across short (i.e. high curvature) axes, leading to oscillations. By averaging gradients across iterations, these oscillations cancel out in the momentum term. However, when the optimizer takes small steps, as is the case in small batch training paired with a small learning rate, parameter updates do not overshoot the minimum across short axes, so there are no oscillations to damp. We illustrate this effect in Appendix˜B through a toy example.
@@ 4.2 Large Batch Training is Sensitive to Hyperparameters
To evaluate robustness to hyperparameters at each batch size, we perform an ablation around the optimal hyperparameters achieving the lowest validation loss for the Adam optimizer. We use the same 30M parameter model trained on the FineWeb-Edu dataset as in Section˜4.1.
We report in Figure˜3 the change in the loss with respect to the lowest value of the loss achieved at every batch size, ranging from $1$ to $4096$. On the x-axis, we report the scaled value of the hyperparameters relative to the value of the hyperparameter that achieved the best loss. We perform this rescaling to make the plots for different batch sizes easier to compare, since the scale of their optimal hyperparameters varies.
Figure 3: Small batch sizes are robust to hyperparameter misspecification. Loss sensitivity to learning rate, $\beta_{1}$, and $\beta_{2}$ using Adam at batch sizes from 1 to 4096. Each curve sweeps a single hyperparameter while holding the others fixed at their optimal values. We observe that smaller batch sizes exhibit broader low-loss regions across all hyperparameters, indicating greater robustness to misspecification. The x-axis shows a scaling of each hyperparameter relative to its optimum value. We parameterize $\beta_{1}$ and $\beta_{2}$ in terms of their decay half-lives, as described in Section˜2. Results are shown for the 30M parameter model trained on 600M tokens of FineWeb-Edu.
@@ 4.3 Adam Hyperparameters: How to Scale Them with Batch Size and Why That Is Necessary
We use our grid search results on the 30M model to test existing scaling laws for Adam hyperparameters and also devise new scaling heuristics for these hyperparameters. Figure˜4 shows the validation loss on FineWeb-Edu for different Adam hyperparameters, namely the learning rate, and decay parameters $\beta_{1}$ and $\beta_{2}$. We similarly plot the loss for different values of the second moment half-life $t_{2}$, which refers to the number of tokens it takes for a gradient’s contribution to the momentum to be reduced to half of its initial value. We provide a more detailed description of $t_{2}$ in Section˜2.
For the learning rate, we observe that the square root scaling recommended by several works and commonly used in practice does not hold (McCandlish et al., 2018; You et al., 2019; Granziol et al., 2022; Malladi et al., 2022; Wang and Aitchison, 2024). In fact, Figure˜4 (left) shows that the learning rate scales more slowly than a square root factor with the batch size. For instance, as we scale the batch size from 1 to 1024, the square root rule would indicate that the corresponding learning rate should be scaled by a factor of 32, whereas we find that a factor of only about 3 empirically works better.
@@
Next, we test our scaling heuristics in a different setting proposed in Xiao (2024). The authors train a decoder-only transformer with 19M non-embedding parameters using Adam on the C4 dataset (Dodge et al., 2021), and only tune the learning rate with $\beta_{1}$ and $\beta_{2}$ fixed to their default values of $0.9$ and $0.95$, respectively. We reproduce their Figure 10(a) results in Figure˜5 (left), where small batch sizes of 16 and 32 perform worse than larger batch sizes of 128 and 256.
Figure 5: Scaling $\beta_{2}$ according to our heuristics significantly improves the performance of small batch sizes. Left: Results from Xiao (2024), showing suboptimal performance for small batch sizes due to fixed Adam hyperparameters. Right: Our results after scaling $\beta_{2}$ such that the decay half-life $t_{2}$ is constant across batch sizes. The smallest batch sizes now perform on par with larger batch sizes.
While our scaling heuristics imply that $\beta_{1}=0.9$ could work well across batch sizes, we recommend scaling $\beta_{2}$ such that $t_{2}$ is fixed. By following this simple one-shot scaling recommendation and without any additional hyperparameter tuning, we achieve significantly different results in Figure˜5 (right). Strikingly, small batch sizes 16 and 32 achieve on par performance with larger batch sizes 128 and 256, therefore implying that: (i) the previous results from Xiao (2024) are largely a consequence of incomplete hyperparameter tuning, (ii) our simple one-shot heuristic generalizes to other settings.
@@ 4.4 SGD Performs Competitively with Adaptive Optimizers at the Billion+ Parameter Scale
Through our experiments with 30M parameter models, we conclude that: (1) small batch sizes can perform on par with or better than larger batch sizes, (2) small batch sizes require less sophisticated optimizers and are more robust to hyperparameter misspecification, (3) in particular, vanilla SGD can be competitive with adaptive optimizers in the small batch regime, and (4) we propose a new scaling heuristic for $\beta_{2}$ such that the half-life $t_{2}$ is fixed, and validate that it leads to improved performance. In Figure˜6, we test these findings at a larger scale by training GPT-2 (124M) (Radford et al., 2019) and GPT-3 (1.3B) (Brown et al., 2020) on the FineWeb dataset (Penedo et al., 2024).
For the larger 1.3B model, hyperparameter tuning of the learning rate for each optimizer was computationally prohibitive; therefore, we used AdamW with the default hyperparameter settings from Brown et al. Brown et al. (2020) (batch size $512,\beta_{1}=0.9,\beta_{2}=0.95,w=0.1$) as our baseline. When using batch size 1 with Adam, we do not perform any hyperparameter tuning – we turn off weight decay and scale the learning rate following our empirical results from Section˜4.3. We compare two scaling heuristics for $\beta_{2}$: keeping $\beta_{2}$ fixed and keeping the half-life $t_{2}$ fixed. We do, however, sweep over four learning rates for Adafactor and SGD, since we have no reference values for these optimizers. Interestingly, we find that SGD performs on par with the AdamW baseline, and Adam and Adafactor with batch size 1 both outperform the AdamW baseline.
@@
Figure 6: Vanilla SGD performs competitively at larger model scales with minimal tuning. We train GPT-3 (1.3B) and GPT-2 (124M) on FineWeb and compare validation loss across optimizers. Left: For GPT-3 (1.3B), vanilla SGD with batch size 1 and no momentum performs on par with the AdamW configuration from Brown et al. (2020), despite having no optimizer state. Our proposed $\beta_{2}$ scaling rule improves Adam’s performance at small batch sizes and outperforms the baseline. Right: After tuning the learning rate of each optimizer on GPT-2 (124M), Adam and Adafactor at batch size 1 match the performance of AdamW at batch size 512, while SGD performs slightly worse.
Figure 7: Minimum memory requirements to train GPT-3 (13B), assuming all parameters and activations are stored in 16-bit precision, gradient checkpointing is used after every transformer block layer, and the backward pass is fused with the optimizer update step (so that the full model gradient never needs to be materialized in memory). The dashed line shows the available memory on an NVIDIA A100 40GB GPU, which is not large enough to store expensive optimizer states or accumulated gradients. We measured the true memory usage of each optimizer by performing fused backward passes in PyTorch on an NVIDIA B200 GPU.
@@ 4.5 Full Fine-tuning on a Budget
In Sections˜4.1, 4.2, 4.3 and 4.4, we studied pretraining across different model scales and discovered that with small batch sizes, optimizers with small or even no state (like Adafactor and SGD) can perform competitively with Adam. We now exploit this observation for memory-efficient fine-tuning.
While pretraining is typically bottlenecked by compute, fine-tuning is instead often bottlenecked by memory. Indeed, fine-tuning a large model on a small dataset does not require many FLOPs, but it does require storing the (large) model and optimizer state in memory. In Figure˜7, we show an example where the GPU has enough memory to store the model, but not enough memory to also store an expensive optimizer like Adam.
@@
A common approach to reduce memory usage during fine-tuning is to freeze the model weights and only train Low-Rank Adaptation (LoRA) (Hu et al., 2022) modules on top of the frozen model, which substantially reduces the number of trainable parameters and also proportionately reduces the size of the optimizer state. However, LoRA often underperforms full parameter fine-tuning (Biderman et al., 2024; Wang and Kanwar, 2023).
We propose to use a small batch size combined with a small optimizer like Adafactor to get the performance benefits of full fine-tuning while having similar memory requirements to LoRA.
Figure 8: Adafactor enables memory-efficient fine-tuning. We fine-tune and evaluate a non-instruction-tuned Gemma 3 (4B) model (Team et al., 2025) on the MATH dataset (Hendrycks et al., 2021). We compare the use of LoRA ($r=16$) against full model fine-tuning, and a small vs. large batch size. For each training method, the width of the plotted bar is proportional to the memory footprint. LoRA uses bfloat16 model weights and float32 adaptors, and Adafactor similarly uses bfloat16 weights but float32 optimizer state. We fix the second moment half-life to be consistent across batch sizes ($\beta_{2}=0.95$ for batch size 16 and $\beta_{2}=0.997$ for batch size 1). We run each optimizer for 5 epochs and sweep over 8 learning rates. We train LoRA using Adam.
We provide an example in Figure˜8 by fine-tuning a non-instruction-tuned Gemma 3 (4B) on the MATH dataset for 5 epochs. As a baseline, we fine-tune the model with Adam and batch size 16 (requiring gradient accumulation), storing all parameters in float32 precision, which requires storing a total of 16 bytes in memory per model parameter. Another baseline is provided by LoRA, which requires only around 2 bytes per model parameter since the pretrained model checkpoint is stored in bfloat16 precision, and the number of trainable parameters for LoRA is negligible compared to the model size.
We make two observations based on our results. First, when using Adam, we can reduce the batch size all the way down to one while following the $\beta_{2}$ scaling heuristic from Equation˜2, eliminating the need to store accumulated gradients in memory. Second, when a small batch size is used, we can replace Adam with Adafactor, which has a much smaller state size. To further reduce the memory use of Adafactor, we store its (tiny) optimizer state in float32, but we store model weights in bfloat16, using stochastic rounding to update the weights after every optimizer step (we discuss the importance of stochastic rounding in Section˜A.3).
@@
Summary: We recommend that practitioners training large models in memory-constrained settings exploit the benefits of small batch sizes (e.g. using optimizers with a small state size) rather than trying to emulate the large batch size setting typically used in industry.
5 Practical Recommendations and Discussion
@@
We show that small batch sizes, even as low as 1, consistently match or outperform large batches across model scales when tuned properly, in both pretraining and fine-tuning. In particular, SGD with batch size 1 performs on par with the AdamW baseline for GPT-3 (1.3B) with minimal tuning. We also propose a simple scaling rule for $\beta_{2}$ based on a fixed second-moment half-life in tokens, which generalizes across model scales. Our results contradict common assumptions about large batch training and gradient accumulation.
Our general recommendation for setting the batch size is to use the smallest batch size that maximizes model throughput (measured in tokens / second) or equivalently the model FLOPs utilization (MFU) Chowdhery et al. (2023). To achieve this objective, the accelerator must spend more time performing computation than moving data between HBM and register memory, which in practice means using a batch size of at least several hundred tokens per device (Austin et al., 2025). When training frontier models on large clusters, this can correspond to a batch size of many millions of tokens. But when training on a single device, the smallest batch size required to achieve high arithmetic intensity could be as low as a few thousand tokens. We therefore recommend avoiding gradient accumulation unless training on multiple devices with multiple model replicas, where the bandwidth between model replicas is a bottleneck. Lastly, when working in a memory-constrained setting, we recommend exploiting the benefits of simple optimizers combined with a small batch size.
@@
Several open questions remain. How do these findings interact with batch size schedules? Can we further reduce weight precision in fine-tuning from 16 to 8 or even 4 bits? Can we design more effective optimizers with a small state, tuned to the small batch size regime? And can we theoretically understand why fixing the second-moment half-life in tokens generalizes across scales?